OpenStudy (anonymous):

identify the graph of the parametric equation: x=sint and y=2cost

OpenStudy (anonymous):

x^2=sin^2(t) y^2/4=cos^2(t) x^2+y^2/4=sin^2+cos^2=1 this is an ellipse. Also we know x=sint,y=cost is a circle, and here we are just stretching along the y axis by a factor of two making an ellipse

OpenStudy (anonymous):

ok. is it possible to graph this on a calculator?

OpenStudy (anonymous):

you can graph it using two functions. one for the top and one for the bottom. solve for y: y^2=4*(1-x^2) y=+sqrt(4*(1-x^2)) y=-sqrt(4*(1-x^2)) on a calculator you need to make two seperate function for the top and bottom half and plot both.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

yw